OpenStudy (anonymous):

A rock dropped from a cliff covers one-third of its total distance to the ground in the last second of its fall. Air resistance is negligible. How high is the cliff?

OpenStudy (anonymous):

distance covered inlast sec. is v+1/2g=v+5=1/3h; v=([2h/g]^1/2 -1)g; v^2=(h/5+1-2[h/5]^1/2)g^2; solve these and eliminatte v you get h

OpenStudy (anonymous):

@ashwinjohn3 what would be the answer in your opinion?

OpenStudy (shamim):

ur answer will b 148.48 m

OpenStudy (shamim):

can i explain my work

OpenStudy (anonymous):

ok use the equation....... Diplacement @ nth second=u+ a(2n-1)/2

OpenStudy (shamim):

no

OpenStudy (shamim):

i wanna use the equation\[h=ut+ \frac{ 1 }{ 2 } g t ^{2}\]

OpenStudy (shamim):

for full height u=0

OpenStudy (shamim):

and my equation become \[h=0+\frac{ 1 }{ 2 } g t ^{2}\]

OpenStudy (shamim):

\[h=\frac{ 1 }{ 2 } \times 10 \times t ^{2}=5 \times t ^{2}\]

OpenStudy (shamim):

its my 1st equation

OpenStudy (shamim):

now i wanna write another equation for 1st 2/3 h height

OpenStudy (shamim):

\[\frac{ 2 }{ 3 }h=ut+\frac{ 1 }{ 2 } \times g t ^{2}=\frac{ 1 }{ 2 } \times 10 \times (t-1)^{2}\]

OpenStudy (shamim):

now u hv to solve these 2 equation to get time t and then u will get the height h

OpenStudy (agent0smith):

You will get two values for t, but one of them will be less than 1 second, making it invalid, as you can see from the question "covers one-third of its total distance to the ground in the *****last second***** of its fall" Solve the two equations shamim gave you, by substituting the first one into the second, then solve for t. The second equation comes from the fact that, if it covers one third of the total distance in the last second, then it must cover the initial 2/3rds of the total distance in (t-1) seconds. \[h = \frac{ 1 }{ 2} g t ^{2}\] \[\frac{ 2 }{ 3} h = \frac{ 1 }{ 2} g (t - 1) ^{2}\]

OpenStudy (anonymous):

Say the height of the cliff is H, time take for fall is T, and g is the acceleration of freefall. \[H=\frac{ 1 }{ 2 }gT ^{2}\] --- (1) \[\frac{ 2 }{3} H=\frac{ 1 }{ 2 }g(T -1)^{2}\] --- (2) Solving you will get 2 answer for T: 0.551 s and 5.45 s But T>1, Hence H = 146 m